# Modern Neural Networks for Small Tabular Datasets: The New Default for Field-Scale Digital Soil Mapping?

### v1

@@ Appendix C Benchmark studies review
Appendix D Hyperparameter search space
D.1 Hyperparameters of artificial neural networks
Table D1: Search spaces for modern ANN architectures. Continuous intervals were sampled uniformly, logarithmic intervals log-uniformly, and width/dimension parameters were sampled from the categorical sets shown.
@@ D.1 Hyperparameters of artificial neural networks
Table D1 summarizes the hyperparameter search spaces of all neural architectures considered in this study, with the exception of TabPFN, which is a foundation model that requires no hyperparameter optimization. Parameters that are shared across similar models are listed under a common name. During hyperparameter optimization, all continuous ranges were sampled uniformly, all the ranges indicated with sub-script "log" were sampled log-uniformly, and all width and dimension parameters were drawn from the categorical sets shown in the table. For the High-Dimensional data group we applied PCA in a joint search, sampling the number of components uniformly from 2–32.
Some model-specific hyperparameters that are not listed in Table D1 were additionally optimized. For AMFormer these were number of prompt tokens $\in\{1,2,4\}$, additive and multiplicative attention tokens $\in\{1,2,4\}$. For ModernNCA these were retrieval parameters, including embedding dimensionality $64\text{--}256$ and soft-NN temperature $0.1\text{--}5.0$ with log-uniform sampling. For TabR these were retrieval parameters, including context size $\{2,4,8,16\}$ and context dropout $0\text{--}0.3$.
D.2 Hyperparameters of classical machine learning algorithms
Table D2: Hyperparameter grids for classical baselines. Bracketed intervals were sampled uniformly; "log" indicates log-uniform sampling.
@@ D.2 Hyperparameters of classical machine learning algorithms
Table D2 lists the search spaces of all classical ML algorithms considered in this study. Sampling rules and the PCA or partial least squares component search (2–32) are identical to those used for the ANNs.
Appendix E Implementation details
@@
All ANN models employed in this study utilize reference implementations of their core architectures. For the model architecture implementations, we use the authors’ reference code provided with the respective publications for ModernNCA (Ye et al., 2025c), TabR (Gorishniy et al., 2024), TabM (Gorishniy et al., 2025), FT-Transformer (Gorishniy et al., 2021), T2G-Former (Yan et al., 2023), AMFormer (Cheng et al., 2024), and ExcelFormer (Chen et al., 2024). For TabPFN (Hollmann et al., 2025), we use the authors’ implementation including the pre-trained model weights. MLP follows the implementation from Gorishniy et al. (2025) and AutoInt follows the PyTorch reimplementation from Gorishniy et al. (2021). RealMLP was reimplemented based on the tabular benchmarking framework from Holzmüller et al. (2024), excluding their custom training pipeline. Numerical feature embeddings (see Section 2.4.2) were based on rtdl-num-embeddings 0.0.12, using an implementation of Quantile-Based Piecewise Linear Encoding that corresponds to the "Q-L" variation from Table 2 in Gorishniy et al. (2022). The benchmarking framework, including our unified model training interface, training loop with early stopping, data preprocessing pipelines (scaling, PCA, numerical embeddings), optimization utilities, cross-validation structure, YAML-based configuration system, and experimental logging, was implemented as part of this work to ensure consistent evaluation across all models.
@@
The evaluation framework was developed in Python 3.10 using PyTorch 2.8.0 (Paszke et al., 2019) for deep learning models, scikit-learn 1.6.1 (Pedregosa et al., 2011) for preprocessing and classical methods, XGBoost 3.0.3 (Chen and Guestrin, 2016) for GBDT implementation, and Optuna 4.4.0 (Akiba et al., 2019) for hyperparameter optimization. The complete implementation, including hyperparameters and experimental results, is openly available (see Code and data availability).

### v2

@@ Appendix E Benchmark studies review
*, †: Ye et al. (2025b), expands on Ye et al. (2025a) with an additional high-dimensional setting (*) and a large-sample setting (†).
Appendix F Hyperparameter search space
F.1 Hyperparameters of artificial neural networks
Table S2: Search spaces for modern ANN architectures. Continuous intervals were sampled uniformly, logarithmic intervals log-uniformly, and width/dimension parameters were sampled from the categorical sets shown.
@@ F.1 Hyperparameters of artificial neural networks
†: All dropout-type hyperparameters (e.g. attn_dropout, ffn_dropout, residual_dropout, context_dropout).
Table S2 summarizes the hyperparameter search spaces of all neural architectures considered in this study, with the exception of TabPFN, which is a foundation model that requires no hyperparameter optimization. Parameters that are shared across similar models are listed under a common name. During hyperparameter optimization, all continuous ranges were sampled uniformly, all the ranges indicated with sub-script "log" were sampled log-uniformly, and all width and dimension parameters were drawn from the categorical sets shown in the table. For the High-Dimensional data group we applied PCA in a joint search, sampling the number of components uniformly from 2–32.
Some model-specific hyperparameters that are not listed in Table S2 were additionally optimized. For AMFormer these were number of prompt tokens $\in\{1,2,4\}$, additive and multiplicative attention tokens $\in\{1,2,4\}$. For ModernNCA these were retrieval parameters, including embedding dimensionality $64\text{--}256$ and soft-NN temperature $0.1\text{--}5.0$ with log-uniform sampling. For TabR these were retrieval parameters, including context size $\{2,4,8,16\}$ and context dropout $0\text{--}0.3$.
F.2 Hyperparameters of classical machine learning algorithms
Table S3: Hyperparameter grids for classical baselines. Bracketed intervals were sampled uniformly; "log" indicates log-uniform sampling.
@@ F.2 Hyperparameters of classical machine learning algorithms
Table S3 lists the search spaces of all classical ML algorithms considered in this study. Sampling rules and the PCA or partial least squares component search (2–32) are identical to those used for the ANNs.
Appendix G Implementation details
@@
All ANN models employed in this study utilize reference implementations of their core architectures. For the model architecture implementations, we use the authors’ reference code provided with the respective publications for ModernNCA (Ye et al., 2025c), TabR (Gorishniy et al., 2024), TabM (Gorishniy et al., 2025), FT-Transformer (Gorishniy et al., 2021), T2G-Former (Yan et al., 2023), AMFormer (Cheng et al., 2024), and ExcelFormer (Chen et al., 2024). For TabPFN (Hollmann et al., 2025), we use the authors’ implementation including the pre-trained model weights. MLP follows the implementation from Gorishniy et al. (2025) and AutoInt follows the PyTorch reimplementation from Gorishniy et al. (2021). RealMLP was reimplemented based on the tabular benchmarking framework from Holzmüller et al. (2024), excluding their custom training pipeline. Numerical feature embeddings were based on rtdl-num-embeddings 0.0.12, using an implementation of Quantile-Based Piecewise Linear Encoding that corresponds to the "Q-L" variation from Table 2 in Gorishniy et al. (2022). The benchmarking framework, including our unified model training interface, training loop with early stopping, data preprocessing pipelines (scaling, PCA, numerical embeddings), optimization utilities, cross-validation structure, YAML-based configuration system, and experimental logging, was implemented as part of this work to ensure consistent evaluation across all models.
@@
The evaluation framework was developed in Python 3.10 using PyTorch 2.8.0 (Paszke et al., 2019) for deep learning models, scikit-learn 1.6.1 (Pedregosa et al., 2011) for preprocessing and classical methods, XGBoost 3.0.3 (Chen and Guestrin, 2016) for GBDT implementation, and Optuna 4.4.0 (Akiba et al., 2019) for hyperparameter optimization. The complete implementation, including hyperparameters and experimental results, is openly available (see Code and data availability in the main manuscript).
Appendix H Glossary
This glossary provides brief and condensed, application-oriented notions of key Artificial Neural Network (ANN) terms as used in this study. For a detailed discussion, we refer to (Bengio et al., 2017; Aggarwal et al., 2018).
Activation function – a (typically) non-linear function applied within a layer that enables the model to (then) capture non-linear relationships.
Architecture – the overall structural design of an ANN, including which layers it contains, how they are connected (e.g., residual/skip connections), and how information flows from inputs to outputs.
Attention – a mechanism that assigns data-dependent importance weights to a set of inputs and forms a weighted combination of them.
Batch – a subset of the data processed together in a single forward and backward pass through the ANN.
Embedding – a dense vector representation of an item in a continuous feature space.
Epoch – one complete pass of all training data through the ANN.
Feed-forward network – an ANN whose computation graph is a directed acyclic graph, with no cycles or feedback connections.
Hyperparameters – variables that control the learning process and are chosen by the practitioner rather than learned during training.
In-context learning – making predictions by conditioning on examples provided at inference time, without updating the ANN’s parameters.
Layer – a building block of an ANN that applies a specific transformation to its input(s).
Learning rate – a hyperparameter used by the optimizer that controls the step size of parameter updates during training.
Loss function – the objective minimized during ANN training.
Multi-head self-attention – multiple self-attention heads computed in parallel on learned projections of the inputs.
Multilayer perceptron – a feed-forward, fully connected neural network composed of stacked linear (affine) layers with non-linear activations; often abbreviated as MLP.
Optimizer – the algorithm that adjusts the ANN’s parameters to minimize the loss function during training.
Parameters – internal variables (e.g., weights and biases) that are learned during training.
Prior-data Fitted Network – a network trained on many synthetic datasets sampled from a prior distribution over supervised learning tasks. At inference time, it approximates Bayesian inference on new datasets in a single forward pass without parameter updates, enabling in-context learning.
Representation learning – the process of automatically learning useful feature representations from raw inputs.
Residual connection – a common form of skip connection that adds a block’s input to its output (element-wise), assuming matching dimensions.
Self-attention – attention applied over elements within the same data sample.
Skip connection – a connection that passes information from an earlier layer directly to a later one, bypassing intermediate layers.
Tokenization – converting raw inputs into discrete units called tokens.
Transformer – an ANN architecture built from stacks of multi-head self-attention and feed-forward blocks, connected with residual connections and normalization, usually with positional encodings to represent order.